OpenStudy (anonymous):

Solve the following inequality.Answer in terms of an improper fraction. 3y + 5 < 10

OpenStudy (anonymous):

\[3y >5\]\[y >5/3\]

OpenStudy (anonymous):

3y + 5 = 10 (subtract 5 from both sides) 3y + 5 - 5 = 10 - 5 (combine like terms) 3y = 10 - 5 3y = 5 (divide by 3 on both sides to get x by itself) 3y/3 = 5/3 x = 5/3

OpenStudy (anonymous):

oops....instead of = sign it is supposed to be >.